# Bochner’s technique in Einstein’s non-symmetric geometry

### v1

@@ Conclusion
The main contribution of this paper (four theorems and ten propositions) is the development of Bochner’s technique for a manifold $(M,G=g+F)$ with an Einstein’s connection. To generalize some geometrical analysis tools for an Einstein’s connection, we assume some additional conditions, for example, (15), (27), (36) and (52), for $F$ and the difference tensor $K$. We define the $f$-curvature operator and the Weitzenböck $f$-curvature operator for NGT spaces, and prove vanishing theorems on the null space of the Bochner and Hodge $f$-Laplacians. We delegate the following for further study: considering NGT spaces associated with special classes of weak metric $f$-manifolds, see [13, 14, 16], and find more applications in NGT and mathematical physics.
Acknowledgments. The author expresses his sincere gratitude to Prof. Milan Zlatanović for numerous fruitful discussions of NGT and its applications. His comments significantly improved both the accuracy and clarity of the manuscript.

### v2

@@ Conclusion
The main contribution of this paper (four theorems and ten propositions) is the development of Bochner’s technique for a manifold $(M,G=g+F)$ with an Einstein’s connection. To generalize some geometrical analysis tools for an Einstein’s connection, we assume some additional conditions, for example, (15) and (36), for $g$, $F$ and the difference tensor $K$, give an example of the Einstein’s connection. We define the $f$-curvature operator and the Weitzenböck $f$-curvature operator for NGT spaces, and prove vanishing theorems on the null space of the Bochner and Hodge $f$-Laplacians. We delegate the following for further study: considering NGT spaces associated with special classes of weak metric $f$-manifolds, see [13, 14, 16], and find more applications in NGT and mathematical physics.
Acknowledgments. The author expresses his sincere gratitude to Professor Milan Zlatanović (University of Niš) for numerous fruitful discussions of NGT and its applications. His comments significantly improved both the accuracy and clarity of the manuscript.